# Double Robust high dimensional alpha test for linear factor pricing model

### v1

@@ A.3 Proof of Theorem 2
where $\delta_{3t}=O_{p}\big{\{}(-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}\|^{2}+\hat{r}^{-2}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\tilde{\boldsymbol{\theta}}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}+\hat{r}^{-2}_{t}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}+\frac{1}{2}\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}-\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}+\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha})^{2}\big{\}}=O_{p}(T^{-1}\log N)$. Then, we have
$\frac{1}{T}\sum_{t=1}^{T}(1-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}+\frac{1}{2}\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}$
$+\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}++\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+\delta_{3t})\hat{{\bf U}}_{t}$
$+\frac{1}{T}\sum_{t=1}^{T}(\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}+\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}-\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha})$
@@
$=$ $\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-1}(1+\delta_{3t}+\delta_{4t})\hat{{\bf D}}^{1/2}\tilde{\boldsymbol{\theta}},$ (8)
@@
$\delta_{4t}=$ $-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\tilde{\boldsymbol{\theta}}$
$+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}+\hat{r}^{-2}_{t}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}2$
$+\frac{1}{2}\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}-\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}$
$+\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}$
@@
$\leq$ $\sqrt{\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-2}(\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha})^{2}}\times\sqrt{\|\frac{1}{T}\sum_{t=1}^{T}\zeta_{1}^{-2}({\bf U}_{t})^{2}\|_{\infty}}$
$=$ $\sqrt{\max_{t}(\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha})^{2}}\sqrt{\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-2}}$
$\leq$ $\sqrt{\max_{t}(\eta_{t}-\eta)^{2}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\alpha}\|^{2}}O_{p}(\zeta_{1})$
$=$ $O_{p}\{T^{-1/2}\log^{1/2}(N)\}.$
In addition,
@@
$T^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}-(1-\eta)\boldsymbol{\alpha})=T^{-1/2}\zeta_{1}^{-1}\sum_{t=1}^{T}{\bf U}_{t}+C_{T}.$ (9)
Similar to the proof of Proposition 2.1, we can derive that $\hat{\zeta}\mathop{\rightarrow}\limits^{p}\zeta.$ Let $u_{N}(y)=y+2\log N-\log\log N$, $T=T^{1/2}||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||_{\infty}\hat{\zeta}^{1/2}$, $T^{c}=T^{1/2}||\hat{{\bf D}}^{-1/2}(\hat{\boldsymbol{\theta}}-(1-\eta)\boldsymbol{\alpha})||_{\infty}\hat{\zeta}^{1/2}$ and recall that $T_{SM}=T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\hat{\zeta}-2\log N+\log\log N$. According to (9), we have
@@
It is clear that, $T\geq T^{1/2}\left\|\hat{\mathbf{D}}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right\|_{\infty}\hat{\zeta}-T^{c}$. Define $q_{1-\alpha}$ to be the $1-\alpha$ quantile of $G(y)$. Combined with Assumption 2 and Lemma 6, we get
$\mathbb{P}\left(T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\hat{\zeta}\geq q_{1-\alpha}+u_{N}(y)\mid H_{1}\right)$
$\geq$ $\mathbb{P}\left(T^{1/2}\left\|\hat{\mathbf{D}}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right\|_{\infty}\cdot\hat{\zeta}-T^{c}\geq\sqrt{q_{1-\alpha}+u_{N}(y)}\mid H_{1}\right)$
$=$ $\mathbb{P}\left(T^{c}\leq T^{1/2}\left\|\hat{\mathbf{D}}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right\|_{\infty}\cdot\hat{\zeta}-\sqrt{q_{1-\alpha}+u_{N}(y)}\mid H_{1}\right)$
$\geq$ $\mathbb{P}\left(T^{c}\leq T^{1/2}\left(\left\|\mathbf{D}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right\|_{\infty}-\left\|\left(\hat{\mathbf{D}}^{-1/2}-\mathbf{D}^{-1/2}\right)(1-\eta)\boldsymbol{\alpha}\right\|_{\infty}\right)\cdot\hat{\zeta}-\sqrt{q_{1-\alpha}+u_{N}(y)}\mid H_{1}\right)$
$\geq$ $\mathbb{P}\left(T^{c}\leq T^{1/2}\left\|\mathbf{D}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right\|_{\infty}\cdot\left(1+O_{p}\left(T^{1/2}\log^{1/2}(NT)\right)\right)\cdot\hat{\zeta}-\sqrt{q_{1-\alpha}+u_{N}(y)}\mid H_{1}\right)\rightarrow 1,$
if $\|\boldsymbol{\alpha}\|_{\infty}\geq\widetilde{C}T^{-1/2}\left\{\log N-2\log\log(1-\alpha)^{-1}\right\}^{1/2}$ for some large enough constant $\widetilde{C}$. The last inequality holds since
$\left\|\left(\hat{\mathbf{D}}^{-1/2}-\mathbf{D}^{-1/2}\right)\boldsymbol{\alpha}\right\|_{\infty}$ $=(1-\eta)\max_{i=1,2,\cdots,N}\frac{\hat{d}_{i}-d_{i}}{\hat{d}_{i}d_{i}}(1-\eta)\alpha_{i}\leq\max_{i=1,2,\cdots,N}\left|1-\frac{d_{i}}{\hat{d}_{i}}\right|\cdot\left\|\mathbf{D}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right\|_{\infty}$
$\leq O_{p}\left(T^{-1/2}\log^{1/2}(NT)\right)\left\|\mathbf{D}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right\|_{\infty}.$

### v2

@@ A.3 Proof of Theorem 2
where $\delta_{3t}=O_{p}\big{\{}(-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}\|^{2}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}-\hat{r}^{-2}_{t}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}-\frac{1}{2}\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}+\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}-\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}-\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha})^{2}\big{\}}=O_{p}(T^{-1}\log N)$. Then, we have
$\frac{1}{T}\sum_{t=1}^{T}(1-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}-\frac{1}{2}\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}$
$-\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}-\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}-\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}+\delta_{3t})\hat{{\bf U}}_{t}$
$+\frac{1}{T}\sum_{t=1}^{T}(\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}+\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}+\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha})$
@@
$=$ $\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-1}(1+\delta_{3t}+\delta_{4t})\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}},$ (6)
@@
$\delta_{4t}=$ $-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{-1/2}\tilde{\boldsymbol{\theta}}$
$+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}-\hat{r}^{-2}_{t}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}2$
$-\frac{1}{2}\hat{r}_{t}^{-2}\|\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|^{2}+\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}$
$-\hat{r}_{t}^{-2}\boldsymbol{V}^{\top}_{t}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}-\hat{r}_{t}^{-2}\tilde{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}(\eta_{t}-\eta)\boldsymbol{\alpha}$
@@
$=$ $\|\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-1}{{\bf U}}_{t}{{\bf U}}_{t}^{\top}{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|_{\infty}\{1+O_{p}(T^{-1/2}\log N)\}$
$=$ $\max_{1\leq i\leq N}\Big{|}\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-1}(\eta_{t}-\eta){U}_{t,i}\sum_{j=1}^{N}{U}_{t,j}\alpha_{j}/d_{j}\Big{|}\{1+O_{p}(T^{-1/2}\log N)\}$
$\leq$ $\max_{1\leq i\leq N}\sqrt{\zeta_{1}^{-2}\frac{1}{T}\sum_{t=1}^{T}{U}_{t,i}^{2}}\sqrt{\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-2}(\eta_{t}-\eta)^{2}\Big{(}\sum_{j=1}^{N}{U}_{t,j}\alpha_{j}/d_{j}\Big{)}^{2}}$
$=$ $O_{p}(1)\sqrt{\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-2}(\eta_{t}-\eta)^{2}\Big{(}\sum_{j=1}^{N}{U}_{t,j}\alpha_{j}/d_{j}\Big{)}^{2}}$
and
$E\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-2}(\eta_{t}-\eta)^{2}\Big{(}\sum_{j=1}^{N}{U}_{t,j}\alpha_{j}/d_{j}\Big{)}^{2}$
$=$ $E\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-2}(\eta_{t}-\eta)^{2}\sum_{j=1}^{N}{U}_{t,j}^{2}\alpha_{j}^{2}/d_{j}^{2}$
$+E\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-2}(\eta_{t}-\eta)^{2}\underset{j_{1}\neq j_{2}}{\sum^{N}\sum^{N}}{U}_{t,j_{1}}\alpha_{j_{1}}/d_{j_{1}}{U}_{t,j_{2}}\alpha_{j_{2}}/d_{j_{2}}$
$=$ $E\frac{1}{T}\sum_{t=1}^{T}\zeta_{2}E(\eta_{t}-\eta)^{2}\sum_{j=1}^{N}\{N^{-1}+O(N^{-1-\delta/2})\}\alpha_{j}^{2}/d_{j}^{2}$
$+E\frac{1}{T}\sum_{t=1}^{T}\zeta_{2}E(\eta_{t}-\eta)^{2}\underset{j_{1}\neq j_{2}}{\sum^{N}\sum^{N}}\{N^{-1}\sigma_{j_{1}j_{2}}+O(N^{-1-\delta/2})\}\alpha_{j_{1}}/d_{j_{1}}\alpha_{j_{2}}/d_{j_{2}}$
$=$ $\frac{1}{T}\sum_{t=1}^{T}\zeta_{2}E(\eta_{t}-\eta)^{2}N^{-1}\boldsymbol{\alpha}^{\top}{{\bf D}}^{-1/2}\mathbf{R}{{\bf D}}^{-1/2}\boldsymbol{\alpha}+O_{p}(N^{-\delta/2}T^{-1}\log N)$
$=$ $O_{p}(N^{-1}T^{-1}\log N\lambda_{\max}(\mathbf{R}))+O_{p}(N^{-\delta/2}T^{-1}\log N)$
$=$ $o_{p}(T^{-1}/\log N),$
where the second inequality holds because of Lemma A.4 in Cheng et al. (2023). Hence, we can obtain that $\|\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}(\eta_{t}-\eta)\boldsymbol{\alpha}\|_{\infty}=o_{p}(\sqrt{T^{-1}/\log N})$. In addition,
@@
$T^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\boldsymbol{\theta}}-\omega\boldsymbol{\alpha})=T^{-1/2}\zeta_{1}^{-1}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){\bf U}_{t}+C_{T},$ (7)
where $\|C_{T}\|_{\infty}=o_{p}(1/\sqrt{\log N})$. Similar to the proof of Proposition 2.1, we can derive that $\hat{\zeta}\mathop{\rightarrow}\limits^{p}\zeta.$ Let $u_{N}(y)=y+2\log N-\log\log N$, $T=T^{1/2}||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||_{\infty}\hat{\zeta}^{1/2}$, $T^{c}=T^{1/2}||\hat{{\bf D}}^{-1/2}(\hat{\boldsymbol{\theta}}-\omega\boldsymbol{\alpha})||_{\infty}\hat{\zeta}^{1/2}$ and recall that $T_{SM}=T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\hat{\zeta}-2\log N+\log\log N$. According to (7), we have
@@
It is clear that, $T\geq T^{1/2}\left\|\hat{\mathbf{D}}^{-1/2}\omega\boldsymbol{\alpha}\right\|_{\infty}\hat{\zeta}-T^{c}$. Define $q_{1-\alpha}$ to be the $1-\alpha$ quantile of $G(y)$. Combined with Assumption 2 and Lemma 6, we get
$P\left(T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\hat{\zeta}\geq q_{1-\alpha}+u_{N}(y)\right)$
$\geq$ $P\left(T^{1/2}\left\|\hat{\mathbf{D}}^{-1/2}\omega\boldsymbol{\alpha}\right\|_{\infty}\cdot\hat{\zeta}-T^{c}\geq\sqrt{q_{1-\alpha}+u_{N}(y)}\right)$
$=$ $P\left(T^{c}\leq T^{1/2}\left\|\hat{\mathbf{D}}^{-1/2}\omega\boldsymbol{\alpha}\right\|_{\infty}\cdot\hat{\zeta}-\sqrt{q_{1-\alpha}+u_{N}(y)}\right)$
$\geq$ $P\left(T^{c}\leq T^{1/2}\left(\left\|\mathbf{D}^{-1/2}\omega\boldsymbol{\alpha}\right\|_{\infty}-\left\|\left(\hat{\mathbf{D}}^{-1/2}-\mathbf{D}^{-1/2}\right)\omega\boldsymbol{\alpha}\right\|_{\infty}\right)\cdot\hat{\zeta}-\sqrt{q_{1-\alpha}+u_{N}(y)}\right)$
$\geq$ $P\left(T^{c}\leq T^{1/2}\left\|\mathbf{D}^{-1/2}\omega\boldsymbol{\alpha}\right\|_{\infty}\cdot\left(1+O_{p}\left(T^{-1/2}\log^{1/2}(N)\right)\right)\cdot\hat{\zeta}-\sqrt{q_{1-\alpha}+u_{N}(y)}\right)\rightarrow 1,$
if $\|\boldsymbol{\alpha}\|_{\infty}\geq\widetilde{C}T^{-1/2}\left\{\log N-2\log\log(1-\alpha)^{-1}\right\}^{1/2}$ for some large enough constant $\widetilde{C}$. Similar to the proof of Lemma 6 in Liu et al. (2024), we can derive that $\max_{1\leq i\leq N}|\hat{d}_{i}^{2}-d_{i}^{2}|=O_{p}(\sqrt{\log N/T})$. Hence, the last inequality holds since
$\left\|\left(\hat{\mathbf{D}}^{-1/2}-\mathbf{D}^{-1/2}\right)\boldsymbol{\alpha}\right\|_{\infty}$ $=\max_{i=1,2,\cdots,N}\frac{\hat{d}_{i}-d_{i}}{\hat{d}_{i}d_{i}}\omega\alpha_{i}\leq\max_{i=1,2,\cdots,N}\left|1-\frac{d_{i}}{\hat{d}_{i}}\right|\cdot\left\|\mathbf{D}^{-1/2}\omega\boldsymbol{\alpha}\right\|_{\infty}$
$\leq O_{p}\left(T^{-1/2}\log^{1/2}(N)\right)\left\|\mathbf{D}^{-1/2}\omega\boldsymbol{\alpha}\right\|_{\infty}.$